(5) Apply Cauchys Integration to EVALUATE the following INTEGRAL: lo x? x4 + 16 dx .
Answer #1


5) १२ 8. x² 2 dx dx 4 I+16 Atio D 22 To Find : dx ~ 14 + 1% Consider de J firos da = ) dy 3+16 + 16 where C is the Semi circ Res [fia), Iz trei J = lim -(vitrai)) fla B-) Strai 2 a Il him 3 Iztuzi (9-152 wai) ) (3-(-Va+v=0)) (8-6-2 Vatic)^ -vai)) (Va2+2;2 -2 (2) i = prai) (-215) (-252 +2532) -ti = - 4v5(-1+0) -si 62 52t2r2i) ll -1-3 882 Hence by Cauchys residue theorem ş